OpenStudy (anonymous):

An observer (O) is located 500 feet from a school (S). The observer notices a bird (B) flying at a 30° angle of elevation from his line of sight. How high is the bird flying over the school? You must show all work to receive full credit.

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

There is a right angled triangle formed tan angle = opposite/adjacent tan30 = x/500 x = 500 *tan(30) = 288.67 feet

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

thnx

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

ok